IR-SPECTROSCOPY OF GLYCINE AND ITS COMPLEXES WITH WATER IN HELIUM NANODROPLETS

\underline{M. LETZNER}, S. A. GRÜN, G. SCHWAAB and M. HAVENITH, Department of Physical Chemistry II, Ruhr-University Bochum, D-44780 Bochum, Germany.

Glycine is the smallest amino acid, and therefore it is of special interest as a model and starting point for theoretical and experimental studies. Whereas the crystalline form of glycine consists of zwitterions NH_{3}^{+} − CH_{2} − COO^{−}^{a}, gas phase glycine is known to exist in the nonionized form NH_{2} − CH_{2} − COOH ^{b}. The interaction between glycine and water has been widely studied using a large variety of theoretical methods. Depending on the theoretical level used, a stabilisation of the zwitterionic form is predicted for complexes containing from 2 to 7 water molecules. In low-temperature Ar matrices a set of characteristic IR absorption bands for the zwitterionic form has been observed. The higher stoichiometry complexes (glycine)⋯(H_{2}O)_{n} with n larger than 3 are demonstrated to be zwitterionic H-bonded complexes ^{c}. The multitude of conformations expected for these glycine-water complexes makes a combination of low temperature and high resolution spectroscopy essential. We want to use the advantages of our experiment to investigate glycine and its complexes with water in helium-nanodroplets at ultracold temperatures in the range from 3000-3800 cm^{−1}. Our measurements were carried out using a high power IR-OPO (cw: 2.7 W) as radiation source and a helium nanodroplet spectrometer. Heliumnanodroplets are formed by expansion of helium at 55 bar through a 5 μm nozzle which is kept at a temperature of 16 K. The status of the project is presented.

aP.-G. Jönsson et al., Acta Crystallogr., Sect. B: Struct. Crystallogr. Cryst. Chem. 28, 1827 (1972)

bG. Junk et al., J. Am. Chem. Soc. 85, 839 (1963)

cR. Ramaekers et al., J. Chem. Phys., 120 (2004)